OpenStudy (isaiahthegreat1):

F is the midpoint of GH. Point G is located at 32 and F is at 105. What is the location of point H? H is located at

OpenStudy (isaiahthegreat1):

this some bs im about to quit this site no one answering anyones questions

OpenStudy (kenrick143):

Wait..

OpenStudy (isaiahthegreat1):

alright my bad.

OpenStudy (kenrick143):

Just arrived haha.

OpenStudy (kenrick143):

Ok. So We have th given

OpenStudy (isaiahthegreat1):

what does that mean?

OpenStudy (kenrick143):

Remember the formula gor midpoint?

OpenStudy (isaiahthegreat1):

the formula for finding the midpoint is (x1 + x2)/2

OpenStudy (isaiahthegreat1):

I really appreciate it :)

OpenStudy (kenrick143):

So we have x=(x1+x2)/2 Substituting.

OpenStudy (kenrick143):

F=(G-H)/2

OpenStudy (isaiahthegreat1):

what is the location of point H?

OpenStudy (kenrick143):

Substituting 105=(H-32)/2 210=H-32 H=242

OpenStudy (isaiahthegreat1):

thank you so much how do i give you medals?

OpenStudy (kenrick143):

I mean the equation is F=(H-G)/2

OpenStudy (kenrick143):

I dont know i just visited here im new.

OpenStudy (isaiahthegreat1):

do you know alot about geometry?

OpenStudy (isaiahthegreat1):

if you do maybe i can pay you through paypal to help me pass my virtual class

OpenStudy (kenrick143):

I just learned but im not a master of geometry. Haha, talking about paying I dont really deserve it. Haha. You could just get a tutor in your country.